OpenStudy (anonymous):

Lim ( x^3 + 5x^2 – 7x – 35) / (x+5) x→-5

OpenStudy (anonymous):

did you check what happens if you plugin \(x=-5\) ?

OpenStudy (anonymous):

thats what im learning now and i think you just do what spacelimbus said just plug it in

OpenStudy (anonymous):

also note that you can write the nominator differently so it will cancel out with the denominator

OpenStudy (anonymous):

btw its numerator

OpenStudy (anonymous):

sorry, I still don't get it. I know the answer is 18 , but how???

OpenStudy (anonymous):

\[x^3 + 5x^2 – 7x – 35=(x+5)(x^2-7) \]

OpenStudy (anonymous):

so you see that the \((x+5)\) from the denominator cancels out, you remain left with \(x^2-7\) now plugin \(x=-5\) will give you \((-5)^2-7=25-7=18\)